OpenStudy (anonymous):

Solve using calculus 1000ln(1+x/400)+600-1.5x-300 Find the maximum

OpenStudy (anonymous):

(266.6677,410.82562)

OpenStudy (anonymous):

can u show working

OpenStudy (anonymous):

i did it on my calculator. sorry don't have time to explain how to do it.

OpenStudy (anonymous):

cmon i can do it on calculator also I need to show working out

OpenStudy (anonymous):

sorry my homework is due in couple of hours. and i need to find out how to do it. Right now i don't get it at all.

OpenStudy (dumbcow):

take derivative and set it equal to 0. solve for x 1+(x/400) = (400+x)/400 d/dx 1000ln((400+x)/400) = 1000/(400+x) => 1000/(400+x) - 1.5 =0 1000 = 1.5x +600 400 = 1.5x 266.67 = x substitute it into original function to find max